OpenStudy (anonymous):

Write the standard equation for the circle center (–6, –8), that passes through (0, 0). (1 point) (x – 6)2 + (y – 8)2 = 10 (x – 6)2 + (y – 8)2 = 196 (x + 6)2 + (y + 8)2 = 14 (x + 6)2 + (y + 8)2 = 100

OpenStudy (anonymous):

\[(x - h)^{2} + (y - k)^{2} = r ^{2}\] plug in the x-value of the center for h and the y-value of the center for k. Then find the radius by finding the distance between (0, 0) and (-6, -8)

OpenStudy (anonymous):

could you answer

OpenStudy (anonymous):

I told you how to do it.

OpenStudy (anonymous):

please

OpenStudy (anonymous):

(x + 6)^2 + (y + 8)^2 = r^2 I did everything except find the radius. Just find the radius and square it.

OpenStudy (anonymous):

I told you how to do it.

OpenStudy (anonymous):

100

OpenStudy (anonymous):

Yes. Now plug that in for r^2 to get (x + 6)^2 + (y + 8)^2 = 100

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome